# Data-Driven Dynamic Factor Modeling via Manifold Learning

### v1

@@ 5 Model Validation
We now turn to an application of our framework to financial stress testing and scenario analysis of equity portfolios. As in [chen1986economic], we consider a standard macroeconomic multi-factor model where stock returns respond to unexpected changes in economy-wide factors. We stress our portfolios using the supervisory scenarios specified in [federal_reserve_2022]. It is important to note that even unconditionally uncorrelated factors can display moderate dependence, for instance, factors can be uncorrelated and yet exhibit significant conditional dependence when a subset of them is learned. As a result the distribution of these factors conditional on an extreme scenario can display strong dependence.
@@
where $\mathbf{Y}_{t}$ is the $n\times 1$ vector of excess returns at time $t$ (where $n$ is the number of assets), $\mathbf{X}_{t}$ is the $p\times 1$ vector of factor realizations at time $t$, $\mathbf{B}=(\mathbf{\beta}_{1},\mathbf{\beta}_{2},\dots,\mathbf{\beta}_{p})$ is the $n\times p$ matrix of factor loadings, where each column $\mathbf{\beta}_{i}=(\beta_{i,1},\beta_{i,2},\dots,\beta_{i,n})^{\top}$ represents the loadings of all assets on factor $i$, and $\mathbf{\varepsilon}_{t}=(\varepsilon_{1,t},\varepsilon_{2,t},\dots,\varepsilon_{n,t})^{\top}$ is the $n\times 1$ vector of idiosyncratic errors. We note that this linear multi-factor model is the special case in our JDKF framework where $G=\mathbf{B}\circ F$ described in Section 3.3, where we recall that under this situation, the diffusion coordinates are computed only on the observations ${x(t)}$ and there is only a lifting operator $\mathbf{H}^{x}$ from the diffusion coordinate space to the space of the observed common factors.
@@ 5.1 Scenario Analysis Models
We consider several benchmark models for conducting the scenario analysis and stress tests. The main premise in scenario analysis models for financial portfolios is that there is a set of risk factors that drive a given asset or portfolio’s value. The goal is then to estimate a given portfolio’s return under different scenarios that involve applying shocks or stresses to some or all of the risk factors. In our macroeconomic multi-factor model described above, for instance, the risk factors are the set of macroeconomic covariates $\{X_{i,t}\}_{i=1}^{p}$. There are several types of scenarios in stress testing models, including hypothetical scenarios chosen by the PM or historical scenarios that have happened throughout financial crises. Figure 4 displays the classic scenario analysis framework and the various types of risk factors, scenarios, and risk metrics that PMs typically deal with in conducting stress tests.
Figure 4: Diagram of the scenario analysis framework.
@@ 5.1.1 Standard Scenario Analysis
Standard Scenario analysis (SSA) for financial portfolios aims to compute a portfolio’s profit and loss (P&L) or overall return under various combinations of stresses to a sub-collection of the risk factors that drive the portfolio’s valuation. It is well known that scenario portfolio numbers calculated under SSA can be quite inaccurate for a wealth of reasons. First, post-stress P&L is computed under the assumption that the un-stressed factors are set to zero; meaning they remain unchanged under stresses to the other factors, however, there is often clear statistical dependence among factors, and in particular conditional dependence. Applying a stress to a sub-collection of factors at a given time period $t$ should also have dynamic propagation effects through future time periods. A good stress test must not only take into account cross-factor dependencies but also the joint dynamics of the common factors themselves to assess how a stress propagates through time considering dynamic factor correlations. Furthermore, scenario analysis often involves dealing with latent (unobserved) factors that may actually be intrinsically driving the dynamics. Finally, one of the chief drawbacks of SSA is its lack of backtesting to validate the accuracy of scenario-conditional P&L, since SSA does not explicitly impose a probability model on the factors, it does not naturally lend itself to backtesting. In [haugh2020scenario], a standard linear Gaussian state space model for the dynamics of the risk factors is proposed and scenario analysis is married with dynamic factor models. A dynamic factor model for scenario analysis has the advantage that it now allows for backtesting the performance of a stress test for a given portfolio. However, when considering equity portfolios where the assets are modeled with a high-dimensional set of common factors and their joint dynamics are highly nonlinear, such a standard state space model no longer applies.
In concrete terms, under the standard multi-factor model for stock returns, if $\mathbf{X_{t}}\in\mathbb{R}^{p}$ is the $p$-dimensional vector of common factors, we define $\mathcal{S}$ to be the set containing the indices of the factors in a scenario (the factors we intend to stress), and thus $\mathcal{S}^{C}$ is the index set of those factors we leave un-stressed and which are not in our scenario. SSA then stresses the components of $\mathbf{X}_{\mathcal{S},t}$ according to a given scenario (e.g. $+20$ on the S&P index, $-2$ on the CPI index, and $+5$ on the US Dollar/Euro exchange rate), and keeps the components in $\mathbf{X}_{\mathcal{S}^{C},t}$ unchanged (i.e. equal to their current value). The new portfolio P&L or overall return $V_{t}$ is then computed with $\mathbf{Y}_{t}$ determined by the scenario and the multi-factor model. By leaving $\mathbf{X}_{\mathcal{S},t}$ unchanged, SSA implicitly assumes that $\mathbb{E}[\mathbf{X}_{\mathcal{S}^{C},t+1}|\mathcal{F}_{t},\mathbf{X}_{\mathcal{S},t+1}]=0$, here $\mathcal{F}_{t}$ denotes a filtration with information up to time $t$.
  1. 1.
  2. 2.
  3. 3.
    Obtain the predicted factor vector under SSA
  4. 4.
We summarize the SSA procedure in the following algorithm, which we will then feed into a historical rolling window backtest in the next section. We let $s$ denote the size of the rolling window. Note that the fitted matrix of factor loadings, $\mathbf{\tilde{B}}$ is treated as an input in this algorithmic format. We denote with $x_{t-s:t}\in\mathbb{R}^{s\times p}$ the matrix of common factors from time $t-s$ up to $t$, and with $y_{t-s:t}\in\mathbb{R}^{s\times n}$ the matrix of asset returns from time $t-s$ up to $t$. We take $x_{\mathrm{actual},\mathcal{S},t+1}$ to be the realized historical scenario.
Algorithm 2 SSA Portfolio Simulation
1:$x_{t-s:t},x_{t+1},\mathbf{\tilde{B}}$ (fitted from $(y_{t-s:t},x_{t-s:t})$)
4:Compute the $1/N$ (or Markowitz) weights $w$
@@ 5.1.2 Static PCA
We employ a PCA-based benchmark to assess the impact of stressed scenarios on asset returns. This benchmark is closer to SSA in that we do not impose model dynamics on the PC components, here, we consider the projection of the vector of factor stresses onto the space spanned by the PC components computed from the factor changes. Unlike SSA however, the un-stressed factors are not assumed to remain unchanged and their co-movement is given from the projection. Given a factor matrix $\mathbf{X}\in\mathbb{R}^{T\times p}$ and a return matrix $\mathbf{Y}\in\mathbb{R}^{T\times n}$, the static PCA procedure comprises a series of steps. First, we estimate the linear multi-factor regression model, then we project the factor changes onto principal components, apply stress scenarios, and predict the corresponding returns from the fitted factor model.
  1. 1.
    Preprocessing and Centering.
    $\Delta\mathbf{X}_{t}=\mathbf{X}_{t}-\mathbf{X}_{t-1}.$
    $\tilde{\mathbf{X}}_{t}=\Delta\mathbf{X}_{t}-\mu_{\Delta\mathbf{X}}.$
  2. 2.
    Estimation of the Multivariate Regression Model.
    We fit the multivariate regression model to obtain $\mathbf{\tilde{B}}$:
    $\mathbf{Y}_{t}=\mathbf{X}_{t}\mathbf{B}+\mathbf{\varepsilon}_{t},$
  3. 3.
    Principal Component Analysis (PCA).
    $\mathbf{X}_{\text{PCA}}=U\Sigma V^{T}.$
    and extract the principal component loading matrix:
    $W=V^{T}.$
  4. 4.
    Applying Stressed Scenarios in PCA Space.
    $\Delta\mathbf{X}=\mathbf{X}_{t+1}^{\text{scenario}}-\mathbf{X}_{t}.$
    Center the perturbation vector:
    ${\Delta\tilde{\mathbf{X}}}=\Delta\mathbf{X}-\mu_{\Delta\mathbf{X}}.$
    Project the stress perturbation onto the PCA component space:
    $\text{proj}_{\text{PCA}}=W(W^{T}{\Delta\tilde{\mathbf{X}}}).$
    Obtain the perturbed scenario factor vector:
  5. 5.
    Prediction of Stressed Returns.
We summarize the static PCA procedure in the following algorithm, which we will feed into our historical rolling window backtest in the next section. We let $s$ denote the size of the rolling window. Note that the fitted matrix of factor loadings, $\mathbf{\tilde{B}}$ is treated as an input in this algorithmic format. We denote with $x_{t-s:t}\in\mathbb{R}^{s\times p}$ the matrix of common factors from time $t-s$ up to $t$, and with $y_{t-s:t}\in\mathbb{R}^{s\times n}$ the matrix of asset returns from time $t-s$ up to $t$. We take $x_{\mathrm{actual},\mathcal{S},t+1}$ to be the realized historical scenario. For notational brevity, in the algorithm below we also assume that the mean of the factor changes, $\mu_{\Delta\mathbf{X}}$, is already added back into $\text{proj}_{\text{PCA}}$
Algorithm 3 Static PCA Portfolio Simulation
1:$x_{t-s:t},x_{t+1},\mathbf{\tilde{B}}$ (fitted from $(y_{t-s:t},x_{t-s:t})$)
2:Apply PCA on the (pre-processed) differenced data $\Delta x_{t-s:t}$
4:Obtain the change vector $\Delta x_{t}=\tilde{x}_{t+1}-x_{t}$
8:Compute the $1/N$ (or Markowitz) weights $w$
@@ 5.1.3 Dynamic PCA
A natural comparison benchmark against method is a dynamic version of Principal Component Analysis (PCA), which provides a simpler alternative to the arguably more complex Diffusion Mapping. Recalling that an empirical covariance matrix $\mathbb{C}$ can be decomposed as $\mathbb{C}=\mathbf{\Gamma}\Lambda\mathbf{\Gamma}^{\top}$, where $\mathbf{\Gamma}$ is the matrix of eigenvectors (columns) and $\Lambda$ is a diagonal matrix of eigenvalues, we define the following PCA-based linear Gaussian state space model as a competing benchmark:
[…]
@@
where $\mathbf{X}_{t}$ is the observed data vector at time $t$; $\mathbf{Z}_{t}$ is the latent state vector (principal component embeddings) at time $t$; $\mathbf{\Gamma}$ is the eigenvector matrix (computed from PCA on the data) serving as the observation (loading) matrix; and $\varepsilon_{t}$ is a Gaussian observation noise with covariance matrix $\mathbf{R}$.
[…]
where $\mathbf{Z}_{t}$ is the latent state vector (principal component embeddings) at time $t$; $\mathbf{A}$ is the transition matrix governing the dynamics of the principal component embeddings; and $\mathbf{Q}$ is the process noise covariance matrix.
  1. 1.
  2. 2.
@@
    From the state space representation we then implement the standard Kalman filter (see Appendix D) to obtain our dynamic estimates $\mathbf{Z}_{t}$ of the PC embeddings. In our empirical experiments we select the dimension of the PC embeddings as $d=5$ such that $>99\%$ of the variance in our data is explained by them.
  3. 3.
    $\mu_{\text{cond, PCA}}=\mathbf{\Gamma}_{\mathcal{S}^{C}}^{+}\mu_{\mathrm{cond}}$
    We obtain $K$ Monte Carlo conditional samples:
    $\{\mathbf{Z}^{(k)}_{t+1}\mid\mathbf{Z}_{t},\mathbf{X}_{\mathcal{S},t}\}_{k=1}^{K}$
  4. 4.
  5. 5.
We summarize the dynamic PCA procedure post-Kalman filter fitting, in the following algorithm, which we will feed into a historical rolling window backtest in the next section. We let $s$ denote the size of the rolling window. Note that the fitted matrix of factor loadings, $\mathbf{\tilde{B}}$, the matrix of eigenvectors $\mathbf{\Gamma}$, and the time $t$ Kalman filter estimate $z_{t}$ for the latent state are all treated as inputs in this algorithmic format. We denote with $x_{t-s:t}\in\mathbb{R}^{s\times p}$ the matrix of common factors from time $t-s$ up to $t$, and $y_{t-s:t}\in\mathbb{R}^{s\times n}$ the matrix of asset returns from time $t-s$ up to $t$.
Algorithm 4 Dynamic PCA Portfolio Simulation
1:$K,z_{t},x_{t+1},\mathbf{\Gamma},\mathbf{\tilde{B}}$ (fitted from $(y_{t-s:t},x_{t-s:t})$)
2:for $k\leftarrow 1$ to $K$ do
3:     Generate $z_{t+1}^{(k)}\mid(z_{t},x_{\mathcal{S},t+1})$
5:     Compute the $1/N$ (or Markowitz) weights $w^{(k)}$
7:end for
5.1.4 JDKF
@@
Unlike SSA, the key advantage of our JDKF framework is that it accounts for dynamic correlations among the factors to compute conditional portfolio return under a given stress scenario, all this without imposing any parametric model on the joint factor dynamics. The conditional sampling procedure described in Section 3.4 allows us to sample dynamic diffusion coordinates $\psi_{t}$ conditional on a scenario in the observation space, $\textbf{X}_{\mathcal{S},t}$. Under the multi-factor model for asset returns, the multivariate regression equation becomes our observation equation, so that we operate under the following state space:
[…]
$\quad\quad\quad\quad\quad\quad\quad\mathbf{X}_{t}=\mathbf{H}_{X}\psi_{t}+v_{X,t},\quad v_{X,t}\sim\mathcal{N}(0,\mathbf{R}_{X})$
@@
where $\mathbf{X}_{t}$ is the observed data vector at time $t$; $\psi_{t}$ is the latent state vector (diffusion coordinates) at time $t$; $\mathbf{H}_{X}$ is the linear lifting operator (computed from applying diffusion maps on the data); and $v_{Y,t}$ is a Gaussian observation noise with covariance matrix $\mathbf{R}_{Y}$, correlated with $v_{X,t}$.
[…]
where $\psi_{t}$ is the latent state vector (diffusion coordinate) at time $t$,; $\mathbf{A}$ is the transition matrix governing the dynamics of the principal component embeddings; and $\mathbf{Q}$ is the process noise covariance matrix.
  1. 1.
  2. 2.
@@
    From the state space equations, we then estimate the JDKF parameters, which include $\mathbf{B}$, $\mathbf{R}_{X}$, $\mathbf{R}_{Y}$, and $\mathbf{R}_{XY}$ by fitting the joint Kalman filter as outlined in Section 3.3 to obtain filtered dynamic diffusion coordinate estimates $\psi_{t}^{\mathrm{JDKF}}$, from which we also obtain $\mathbf{X}_{t}^{\mathrm{JDKF}}=\mathbf{H}_{X}\psi_{t}^{\mathrm{JDKF}}$
  3. 3.
    $\mu_{\mathrm{cond,diff}}=\mathbf{H}_{\mathcal{S}^{C}}^{+}\mu_{\mathrm{cond}}$
    We obtain $K$ Monte Carlo conditional samples:
    $\{\psi^{(k)}_{t+1}\mid\psi_{t},\mathbf{X}_{\mathcal{S},t}\}_{k=1}^{K}$
  4. 4.
    $\mathbf{Y}_{t+1,\mathrm{JDKF}}^{(k)}=\mathbf{B}\mathbf{H}_{X}\psi_{t+1}^{(k)}$
  5. 5.
We summarize the JDKF procedure post-Kalman filter fitting, in the following algorithm, which we will feed into a historical rolling window backtest in the next section. We let $s$ denote the size of the rolling window. Note that the fitted matrix of factor loadings, $\mathbf{B}$, the linear lifting operator $\mathbf{H}_{X}$, and the time $t$ Kalman filter estimate $\psi_{t}$ for the latent state are all treated as inputs to the algorithm. We denote with $x_{t-s:t}\in\mathbb{R}^{s\times p}$ the matrix of common factors from time $t-s$ up to $t$, and $y_{t-s:t}\in\mathbb{R}^{s\times n}$ the matrix of asset returns from time $t-s$ up to $t$.
Algorithm 5 JDKF Portfolio Simulation
$K,\hat{\psi}_{t},\mathbf{B},\mathbf{H}_{x},x_{t+1}$
2:for $k\leftarrow 1$ to $K$ do
3:     Generate $\hat{\psi}_{t+1}^{(k)}\mid(\hat{\psi}_{t},x_{\mathcal{S},t+1})$
5:     Compute the $1/N$ (or Markowitz) weights $w^{(k)}$
7:end for

### v2

@@ 6 Application to Financial Stress Testing
We now turn to an application of our framework to financial stress testing and scenario analysis of equity portfolios. As in [11], we consider a standard macroeconomic multi-factor model where stock returns respond to unexpected changes in economy-wide factors. We stress our portfolios using the supervisory scenarios specified in [5]. It is important to note that even unconditionally uncorrelated factors can display moderate dependence, for instance, factors can be uncorrelated and yet exhibit significant conditional dependence when a subset of them is learned. As a result the distribution of these factors conditional on an extreme scenario can display strong dependence.
@@
where $\mathbf{Y}_{t}$ is the $n\times 1$ vector of excess returns at time $t$ (where $n$ is the number of assets), $\mathbf{X}_{t}$ is the $p\times 1$ vector of factor realizations at time $t$, $\mathbf{B}=(\mathbf{\beta}_{1},\mathbf{\beta}_{2},\dots,\mathbf{\beta}_{p})$ is the $n\times p$ matrix of factor loadings, where each column $\mathbf{\beta}_{i}=(\beta_{i,1},\beta_{i,2},\dots,\beta_{i,n})^{\top}$ represents the loadings of all assets on factor $i$, and $\mathbf{\varepsilon}_{t}=(\varepsilon_{1,t},\varepsilon_{2,t},\dots,\varepsilon_{n,t})^{\top}$ is the $n\times 1$ vector of idiosyncratic errors. We note that this linear multi-factor model is the special case in our JDKF framework where $G=\mathbf{B}\circ F$ described in Section 5, where we recall that under this situation, the diffusion coordinates are computed only on the observations x(t){x(t)} and there is only a lifting operator $\mathbf{H}^{x}$ from the diffusion coordinate space to the space of the observed common factors.
@@ 6.1 Scenario Analysis Models
We compare JDKF against three benchmark approaches for stress testing equity portfolios under macroeconomic scenarios. Each method addresses the challenge of predicting portfolio returns conditional on shocks to a subset $\mathcal{S}$ of risk factors, while accounting for co-movements in the remaining factors $\mathcal{S}^{C}$. Figure 4 provides an overview of the scenario analysis framework common to all methods.
Figure 4: Diagram for scenario analysis in stress testing. Input factors (macroeconomic variables, market data, portfolio composition) feed into scenario definitions, which can be historical (e.g., 2008 crisis, COVID-19), hypothetical (e.g., interest rate shock, oil price surge), or generated via Monte Carlo simulations. We then assess the portfolio change under stress and evaluate it through risk metrics such as Value-at-Risk (VaR), Expected Shortfall, stress losses, and regulatory capital requirements (Basel, CCAR).
@@ 6.1.1 Standard Scenario Analysis
Standard Scenario Analysis (SSA) for financial portfolios aims to compute a portfolio’s profit and loss (P&L) or overall return under various combinations of stresses to a sub-collection of the risk factors that drive the portfolio’s valuation. It is well known that scenario portfolio numbers calculated under SSA can be quite inaccurate for several reasons. First, post-stress P&L is computed under the assumption that the un-stressed factors remain unchanged, however, there is often clear statistical dependence among factors, and in particular conditional dependence. Applying a stress to a sub-collection of factors at a given time period $t$ should also have dynamic propagation effects through future time periods. A good stress test must not only take into account cross-factor dependencies but also the joint dynamics of the common factors themselves to assess how a stress propagates through time.
One drawback of SSA is its lack of backtesting to validate the accuracy of scenario-conditional P&L. Since SSA does not explicitly impose a probability model on the factors, it does not naturally lend itself to backtesting. In [22], they propose a standard linear Gaussian state space model for the dynamics of the risk factors and scenario analysis is married with dynamic factor models. A dynamic factor model for scenario analysis has the advantage that it now allows for backtesting the performance of a stress test for a given portfolio. However, when considering equity portfolios where the assets are modeled with a high-dimensional set of common factors and their joint dynamics are highly nonlinear, such a standard state space model no longer applies.
Under the standard multi-factor model, SSA stresses the components of $\mathbf{X}_{\mathcal{S},t}$ according to the given scenario while keeping components in $\mathbf{X}_{\mathcal{S}^{C},t}$ unchanged (i.e., equal to their current value). The portfolio return is then computed as $V_{t}^{\text{SSA}}=\mathbf{w}\cdot\mathbf{Y}_{t}^{\text{SSA}}$, where $\mathbf{Y}_{t}^{\text{SSA}}=\tilde{\mathbf{B}}\mathbf{X}_{t}^{\text{SSA}}$ and $\tilde{\mathbf{B}}$ is estimated via OLS regression. We provide the complete algorithmic details in Appendix F.1. This makes the implicitly assumption $\mathbb{E}[\mathbf{X}_{\mathcal{S}^{C},t+1}|\mathcal{F}_{t},\mathbf{X}_{\mathcal{S},t+1}]=0$, which is a fundamental limitation. It ignores dynamic correlations: stressing inflation and unemployment should affect expectations of unstressed risk factors, yet SSA holds the unstressed factors fixed. Our JDKF framework addresses this by learning the conditional distribution of unstressed factors given stressed ones through the joint covariation structure captured in diffusion coordinates.
@@ 6.1.2 Static PCA
We employ a PCA-based benchmark to assess the impact of stressed scenarios on asset returns. This benchmark is closer to SSA in that we do not impose model dynamics on the PC components; however, unlike SSA, the un-stressed factors are not assumed to remain unchanged and their co-movement is given from the projection onto the principal component space.
Given a factor matrix $\mathbf{X}\in\mathbb{R}^{T\times p}$ and a return matrix $\mathbf{Y}\in\mathbb{R}^{T\times n}$, we first estimate the linear multi-factor regression model to obtain $\tilde{\mathbf{B}}$, then compute principal components from the differenced and centered factor data $\tilde{\mathbf{X}}_{t}=\Delta\mathbf{X}_{t}-\boldsymbol{\mu}_{\Delta X}$. For each specified PCA dimension $d$, we compute the PCA transformation and extract the principal component loading matrix $\mathbf{W}=\mathbf{V}^{T}$ from the singular value decomposition.
[…]
For each backtest, we select the PCA dimension $d$ that minimizes mean absolute error (MAE), giving Static PCA the best possible performance. Complete implementation details including the preprocessing steps are provided in Algorithm 3 (Appendix F.2).
@@ 6.1.3 Dynamic PCA
A natural comparison benchmark is a dynamic version of Principal Component Analysis (PCA), which provides a simpler alternative to the arguably more complex diffusion mapping. Dynamic PCA combines principal component dimensionality reduction with a linear-Gaussian state-space model for the PC time series. The model structure mirrors JDKF but replaces diffusion coordinates with PCs.
We define the following PCA-based linear Gaussian state space model:
Observation Equation: $\mathbf{X}_{t}=\boldsymbol{\Gamma}\mathbf{Z}_{t}+\boldsymbol{\varepsilon}_{t},\quad\boldsymbol{\varepsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{R})$ (6.1)
State Equation: $\mathbf{Z}_{t}=\mathbf{A}\mathbf{Z}_{t-1}+\boldsymbol{\eta}_{t},\quad\boldsymbol{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{Q})$ (6.2)
@@
where $\mathbf{X}_{t}$ is the observed data vector at time $t$; $\mathbf{Z}_{t}$ is the latent state vector (principal component embeddings) at time $t$; $\boldsymbol{\Gamma}$ is the eigenvector matrix (computed from PCA on the data) serving as the observation (loading) matrix; $\mathbf{A}$ is the transition matrix governing the dynamics of the principal component embeddings; and $\mathbf{Q}$ is the process noise covariance matrix.
@@
We estimate the transition matrix $\mathbf{A}$ by fitting a VAR(1) process to the PC components computed from the data. We then implement the standard Kalman filter (see Appendix E) to obtain dynamic estimates $\hat{\mathbf{Z}}_{t}$ of the PC embeddings. In our empirical experiments we select the dimension of the PC embeddings as $d=5$ such that $99\%$ of the variance in our data is explained.
Once we have obtained our estimates for $\mathbf{Z}_{t}$, we perform conditional sampling given a scenario to obtain the stressed factor vector. The conditional sampling procedure closely follows that outlined in Section 5.4 with the exception that we replace the lifting operator matrix with $\boldsymbol{\Gamma}$ and the diffusion coordinates with $\mathbf{Z}_{t}$. We obtain $K$ Monte Carlo conditional samples $\{\mathbf{Z}^{(k)}_{t+1}|\mathbf{Z}_{t},\mathbf{X}_{\mathcal{S},t}\}_{k=1}^{K}$, predict post-stress asset returns for each sample as $\mathbf{Y}^{(k)}_{t+1,\text{Dynamic PCA}}=\tilde{\mathbf{B}}\boldsymbol{\Gamma}^{T}\mathbf{Z}^{(k)}_{t+1}$, and obtain the final dynamic PCA post-stress returns by averaging across all Monte Carlo predictions.
The complete algorithmic details are provided in Appendix F.3, and summarized in Algorithm 4.
6.1.4 Linear Multi-Factor JDKF
@@
Unlike SSA, the key advantage of our JDKF framework is that it accounts for dynamic correlations among the factors to compute conditional portfolio return under a given stress scenario, all this without imposing any parametric model on the joint factor dynamics. The conditional sampling procedure described in Section 5.4 allows us to sample dynamic diffusion coordinates $\boldsymbol{\psi}_{t}$ conditional on a scenario in the observation space, $\mathbf{X}_{\mathcal{S},t}$.
Under the multi-factor model for asset returns, the multivariate regression equation becomes our observation equation, so that we operate under the following state space:
Observation Equations: $\mathbf{Y}_{t}=\mathbf{B}\mathbf{X}_{t}+\boldsymbol{\varepsilon}_{t}=\mathbf{B}\mathbf{H}^{X}\boldsymbol{\psi}_{t}+\mathbf{v}_{Y,t},\quad\mathbf{v}_{Y,t}\sim\mathcal{N}(\mathbf{0},\mathbf{R}_{Y})$
$\mathbf{X}_{t}=\mathbf{H}^{X}\boldsymbol{\psi}_{t}+\mathbf{v}_{X,t},\quad\mathbf{v}_{X,t}\sim\mathcal{N}(\mathbf{0},\mathbf{R}_{X})$ (6.3)
State Equation: $\boldsymbol{\psi}_{t+1}=\mathbf{A}\boldsymbol{\psi}_{t}+\boldsymbol{\eta}_{t},\quad\boldsymbol{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{Q})$
@@
where $\mathbf{X}_{t}$ is the observed data vector at time $t$; $\boldsymbol{\psi}_{t}$ is the latent state vector (diffusion coordinates) at time $t$; $\mathbf{H}^{X}$ is the linear lifting operator (computed from applying diffusion maps on the data); $\mathbf{v}_{Y,t}$ is a Gaussian observation noise with covariance matrix $\mathbf{R}_{Y}$, correlated with $\mathbf{v}_{X,t}$; $\mathbf{A}$ is the transition matrix governing the dynamics of the diffusion coordinates; and $\mathbf{Q}$ is the process noise covariance matrix.
@@
We first apply diffusion maps on the common factors data matrix $\mathbf{X}\in\mathbb{R}^{T\times p}$ to obtain the matrix of diffusion coordinates $\boldsymbol{\psi}\in\mathbb{R}^{T\times\ell}$, the lifting operator $\mathbf{H}^{X}$, and the transition matrix $\mathbf{A}=(\mathbf{I}+\boldsymbol{\Lambda})$. We then estimate the JDKF parameters $\mathbf{B}$, $\mathbf{R}_{X}$, $\mathbf{R}_{Y}$, and $\mathbf{R}_{XY}$ by fitting the joint Kalman filter as outlined in Section 5 to obtain filtered dynamic diffusion coordinate estimates $\boldsymbol{\psi}^{\text{JDKF}}_{t}$.
Using the estimated diffusion coordinates from JDKF, we perform conditional sampling given the scenario in the factor observation space following the procedure outlined in Section 5.4. We obtain $K$ Monte Carlo conditional samples $\{\boldsymbol{\psi}^{(k)}_{t+1}|\boldsymbol{\psi}_{t},\mathbf{X}_{\mathcal{S},t}\}_{k=1}^{K}$, predict post-stress asset returns for each conditional Monte Carlo sample as $\mathbf{Y}^{(k)}_{t+1,\text{JDKF}}=\mathbf{B}\mathbf{H}^{X}\boldsymbol{\psi}^{(k)}_{t+1}$, and obtain the JDKF post-stress returns by averaging across all Monte Carlo predictions: $\mathbf{Y}_{t+1,\text{JDKF}}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{Y}^{(k)}_{t+1,\text{JDKF}}$.
The complete implementation details are provided in Appendix F.4, including the step-by-step procedure described in Algorithm 5.